OpenStudy (anonymous):

Select the GCF of these numbers. 48 and 60

OpenStudy (anonymous):

@zepdrix i did this one ......i think 2^2x3

OpenStudy (anonymous):

x is multiplication symbol

OpenStudy (ineedhelplz):

2 x 2 x 3 = 12

OpenStudy (anonymous):

mindblown

OpenStudy (anonymous):

hahhaa

OpenStudy (ineedhelplz):

?

OpenStudy (anonymous):

?......am i not correctt?

OpenStudy (♪chibiterasu):

You should use * for multiplication. Using x will make us confuse it with variables.

OpenStudy (ineedhelplz):

Yes you are

OpenStudy (♪chibiterasu):

Also 2 x 2 x 3 can be simplified to 12, so yes

OpenStudy (ineedhelplz):

Sorry @♪Chibiterasu

OpenStudy (♪chibiterasu):

I mean context makes it understandable, but js

OpenStudy (ineedhelplz):

He was using x so I assumed he would get confused if I used *

OpenStudy (anonymous):

ill show you what the actual choices are

OpenStudy (anonymous):

Select the GCF of these numbers. 48 and 60 2^2 ·3 2· 11^2 3^2 2^3 · 5 13· 19^3 ·23^2

OpenStudy (anonymous):

so a

OpenStudy (ineedhelplz):

Yes

OpenStudy (♪chibiterasu):

is the ^ supposed to be an exponent?

OpenStudy (anonymous):

yes sir

jhonyy9 (jhonyy9):

48 2 24 2 12 2 6 2 3 3 1 60 3 20 2 10 2 5 5 1 what is the great common factor ? hope this will help you

OpenStudy (♪chibiterasu):

You didn't even list out all the factors. 12 is also a factor of 60.

OpenStudy (♪chibiterasu):

Also it's still a. 2^2 = 2*2 = 4 4*3 = 12

OpenStudy (anonymous):

thanks for all the help